OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -20, -16, -12, -8, .

OpenStudy (anonymous):

it is add 4

OpenStudy (anonymous):

-20, (-4) -16, (4) -12, (4) -8, (4) -4...

OpenStudy (mathstudent55):

The first term is \(a_1 = -20\) Each next term is 4 more than the previous term. Term Number Calculation Term 1 -20 = -20 -20 2 -20 + 4 =-20 + 4 * 1 -16 3 -20 + 4 + 4 = -20 + 4 * 2 -12

OpenStudy (anonymous):

the 9th term will be positive 12

OpenStudy (mathstudent55):

Low look at the pattern in the calculation column, right side.

OpenStudy (anonymous):

@Lethal sp would it be an = -20 + 4(n + 1)

OpenStudy (mathstudent55):

n = 1 \(a_n = a_1 = -20\) n = 2 \(a_n = a_2 = a_1 + 4(n - 1) = -20 + 4 = -16\) n = 3 \(a_n = a_3 = a_1 + 4(n - 1) = -20 + 8 = -12\) The nth term is \(a_n = -20 + 4(n - 1) \)

OpenStudy (agent0smith):

@melacho a minor correction an = -20 + 4(n - 1)

OpenStudy (anonymous):

THANK YOU

OpenStudy (wolf1728):

Couldn't the equation be written simpler such as: an = -20 + n*4 ? n = 1 -20 + 1*4 = -16 n = 2 -20 + 2*4 = -12

OpenStudy (agent0smith):

You could simplify it to an = -24+ 4n

OpenStudy (wolf1728):

So then a0 = -24 and a1 = -20?